OpenStudy (anonymous):

***I only need the second question*** During World War I, the Germans had a gun called Big Bertha that was used to shell Paris. The shell had an initial speed of 2.41 km/s at an initial inclination of 61.5 to the horizontal. The acceleration of gravity is 9.8 m/s2 . How far away did the shell hit? Answer in units of km How long was it in the air? Answer in units of s

OpenStudy (***[isuru]***):

hi,  hence u r confident with the 1st one I think u don't need much of an explanation for the 2nd one before all let's just convert Km/s into m/s 2.41 km/s = 2410m/s To solve the second part First we have to find Vy Vy = 2410 m/s * sin 61.5 = 2410 m/s * 0.8788 =2117.908 m/s When the shell reaches it's maximum height the vertical velocity component will be zero based on that theorem let's apply linear motion equation V = u + at 0 = 2117.908 m/s - 9.8 m/S^2 * (t) t = 216.1130 s As the shell has to go upward and comes back to earth the shell will require a same amount of time in order to reach earth again so the total time that shell will spend on the air will be 2*t = 432.226 s I think this should be the answer (though they are in decimal values ) hope this will help ya!!!